ok.
i have this problem
ln(x-2) + ln(2x-3) = 2lnx
please help me
i have no understanding of this problem

Answer by nerdybill(5404)

ln(x-2) + ln(2x-3) = 2lnx
ln(x-2)(2x-3) = 2ln x
ln(x-2)(2x-3) = ln x^2
(x-2)(2x-3) = x^2
2x^2-3x-4x+6 = x^2
2x^2-7x+6 = x^2
x^2-7x+6 = 0
(x-6)(x-1) = 0
x = {6,1}